OpenStudy (anonymous):

3x2 < (3x)2 For what value of x is the statement above FALSE? (A) -3 (B) 0 (C) 1/3 (D)1 (E) For no value of x

OpenStudy (zarkon):

either solve it or just plug in the values they gave you and see if any don't work

OpenStudy (anonymous):

3x\(^2\) = 3*x*x (3x)\(^2\) = (3x)(3x) = 3 * 3 * x * x = 9x\(^2\) Which do you think grows faster? :-)

OpenStudy (anonymous):

um what

OpenStudy (anonymous):

what does that mean

OpenStudy (zarkon):

\[3x^2<(3x)^2=9x^2\] \[0<9x^2-3x^2=6x^2\] \[0<6x^2\]...what \(x\)'s make this false?